OpenStudy (anonymous):

Simplify the square root of -72. (Choices in comments.)

OpenStudy (anonymous):

\[-6\sqrt{2}\] \[6\sqrt{-2}\] \[6\sqrt{2i}\] \[6i \sqrt{2}\]

OpenStudy (anonymous):

I believe it is \[6i \sqrt{2}\]

OpenStudy (lochana):

hint: \[i^2 = -1\]

OpenStudy (lochana):

yes. you are correct

OpenStudy (anonymous):

Thanks.

OpenStudy (lochana):

you are welcome:)